# Emergence of Superposition: Unveiling the Training Dynamics of Chain of Continuous Thought

### v2

@@ Abstract
Previous work shows that the chain of continuous thought (continuous CoT) improves the reasoning capability of large language models (LLMs) by enabling implicit parallel thinking, and a subsequent work provided theoretical insight by showing that a two-layer transformer equipped with continuous CoT can efficiently solve directed graph reachability by maintaining a superposition of multiple reasoning traces in the continuous thought. However, it remains unclear how the superposition mechanism is naturally learned from gradient-based training methods. To fill this gap, we theoretically analyze the training dynamics of a simplified two-layer transformer on the directed graph reachability problem to unveil how the superposition mechanism emerges during training in two training stages – (i) a thought-generation stage that autoregressively expands the continuous thought, and (ii) a prediction stage that converts the thought into the final answer. Our analysis reveals that during training using continuous thought, the index-matching logit, an important quantity which reflects the strength of the model’s local search ability, will first increase and then remain bounded under mild assumptions. The bounded index-matching logit effectively balances exploration and exploitation during the reasoning process: the model will exploit local problem structures to identify plausible search traces, and assign comparable weights to multiple such traces to explore when it is uncertain about which solution is correct, which results in superposition. Our experimental results tracking the growth of logits further validate our theory. Our code is available at https://github.com/Ber666/emergence-of-superposition.
@@ 1 Introductions
Importantly, our analysis of the thought generation training stage reveals why the superposition can emerge even if the training data only presents one demonstration for each training sample. Our theoretical analysis as well as experimental results show that when training with continuous thought (i.e., the Coconut training method in Section 3 and Section 5), the index-matching logit, an important quantity that measures the strength of model’s local search capability, will remain bounded under mild conditions, which is in contrast to many previous analysis on transformer logit without continuous thought (e.g., Tian et al. (2023a); Nichani et al. (2024a); Nguyen & Nguyen-Tang (2025)) where the logit will grow logarithmically and thus unbounded. A bounded index-matching logit can balance exploration and exploitation: if the logit is too small, the model cannot even perform local search, resulting in a nearly random guess in the next step; if the logit is too large, the model might over-confidently commit to one of the plausible search traces merely depending on local features (e.g., the indegree of a node in our case) even if it is uncertain about the solution, and thus early discard the correct path. A bounded index-matching logit encourages the model to exploit the local structure while explore multiple plausible solutions by assigning comparable weights to them, which naturally results in superposition. This answers the question of Zhu et al. (2025) why superposition can emerge during training.
@@ 1.1 Related works
Chain-of-thought (CoT) (Wei et al., 2022) is a simple yet effective test time scaling method to enhance LLM’s reasoning capability. It can either be prompt-based only (Khot et al., 2022; Zhou et al., 2022) or be included in the training sample to create high-quality training data (Yue et al., 2023; Yu et al., 2023; Wang et al., 2023a; Shao et al., 2024). Beyond empirical study, many theoretical works also explore the advantages of the CoT method. For example, Liu et al. (2022); Feng et al. (2023); Merrill & Sabharwal (2023); Li et al. (2024b) shows that CoT can improve the expressivity of transformers. Zhu et al. (2024) studies the importance of CoT for two-hop reasoning via training dynamics. Wen et al. (2024); Kim & Suzuki (2024) studies how CoT in the training data can improve the sample efficiency of transformers. Instead of the text-based CoT, this paper studies continuous CoT where the “thinking tokens” lie in a latent continuous space and do not need to be converted to discrete tokens.
@@
A recent line of work studies latent space reasoning, a novel paradigm beyond text-based CoT (Goyal et al., 2023; Wang et al., 2023b; Pfau et al., 2024; Su et al., 2025; Hao et al., 2024). For example, Goyal et al. (2023) proposed to use pause tokens, which are learnable tokens that are inserted into the original text to increase the computation space. Later, London & Kanade (2025) theoretically shows that the pause token can strictly increase the expressivity of the transformer. Similarly, Pfau et al. (2024) studies filler tokens, which also increase the computation space of LLMs. Wang et al. (2023b) proposed to use planning tokens at the beginning of the response generation to improve the reasoning capability. Su et al. (2025) proposed to use abstract tokens in a latent space to enhance the reasoning performance while reducing the inference cost. The most related work is Hao et al. (2024), which proposes to use continuous CoT for reasoning. A follow-up work Zhu et al. (2025) theoretically shows the advantage of continuous CoT via expressivity. Our work takes a further step by analyzing the training dynamics of continuous CoT.
@@
There is a rich line of literature studying the optimization of transformer-based models (Jelassi et al., 2022; Bietti et al., 2023; Mahankali et al., 2023; Fu et al., 2023; Tian et al., 2023a; b; Zhang et al., 2024; Li et al., 2024a; Huang et al., 2024; Guo et al., 2024). A line of more recent works focus on the understanding of reasoning abilities or patterns of transformers via training dynamics, including the induction heads (Nichani et al., 2024a), the reversal curse (Zhu et al., 2024), CoT (Wen et al., 2024; Kim & Suzuki, 2024), factual recall (Nichani et al., 2024b), in context two hop reasoning (Guo et al., 2025), out of context reasoning (Huang et al., 2025), etc. Along the line, our paper aims to understand the internal mechanism of continuous CoT and why superposition emerges via the analysis of training dynamics.
@@ 3 Analysis of the Thought Generation Stage
Zhu et al. (2025) constructs a solution for a two-layer transformer, where the first layer mainly performs copy (e.g., the $i$-th special edge token <e> will copy the information of its corresponding source node $\texttt{s}_{i}$ and target node $\texttt{t}_{i}$). Since the copy mechanism has been widely studied (Nguyen & Nguyen-Tang, 2025), as well as its formation via training dynamics (Nichani et al., 2024a), we mainly focus on the dynamics after the copy mechanism has been established. Thus, we analyze the dynamics of the second layer of the transformer.
@@
where $\mathsf{E}_{\texttt{s}}(v)\in\mathbb{R}^{d}$ and $\mathsf{E}_{\texttt{t}}(v)\in\mathbb{R}^{d}$ map token $v\in\mathsf{Voc}$ to different subspaces of $\mathbb{R}^{d}$. For example, as in the construction of Zhu et al. (2025), we can set $d=3M$, and $\mathsf{E}_{\texttt{s}}(\cdot)$, $\mathsf{E}_{\texttt{t}}(\cdot)$ and $\mathsf{E}(\cdot)$ each corresponds to $M$ different non-zero entries. This is also similar to previous work Chen et al. (2025); Nguyen & Nguyen-Tang (2025) where $\mathsf{E}_{\texttt{s}}(\cdot)$ and $\mathsf{E}_{\texttt{t}}(\cdot)$ can be viewed as previous token heads. We make the following assumptions on the embedding $\mathsf{E}_{\texttt{s}}(\cdot)$, $\mathsf{E}_{\texttt{t}}(\cdot)$ and $\mathsf{E}(\cdot)$:
@@
where $\mathbf{V},\mathbf{W}\in\mathbb{R}^{d\times d}$ are attention parameters and $\sigma(\cdot):\mathbb{R}\to\mathbb{R}$ is an activation function that determines the range of attention scores, and $\bm{\xi}=(\xi_{v})_{v\in\mathsf{Voc}}\in\mathbb{R}^{M}$ is the output logit vector for each token in the vocabulary. Similar to the analysis in Nguyen & Nguyen-Tang (2025), we adopt the linear attention $\sigma({\bm{h}}^{\top}\mathbf{W}{\bm{h}}_{i})={\bm{h}}^{\top}\mathbf{W}{\bm{h}}_{i}$, fix $\mathbf{V}={\bm{I}}$ and use the index-matching reparameterization
@@
The first loss $\mathcal{L}^{\mathrm{BFS}}$ explicitly encourages the model to predict any nodes in $\mathcal{N}_{C+1}$. However, in practice, it is costly and even impossible to search over the entire solution space exhaustively; instead, we usually present only one demonstration for each task instance during training (in our setting, only one path $p$ per instance $(\mathcal{G},\texttt{r},\texttt{c}_{1},\texttt{c}_{2})$), which corresponds to the second loss $\mathcal{L}^{\mathrm{coco}}$ and aligns with the practical setting where chain of thought data can be used for supervision.
@@
By Theorem 1, as long as $F(0)<d_{p_{c+1}}<d_{\max}$, we have $\mu(t)\to\mu^{*}>0$. Compared to many previous work (Tian et al., 2023a; Nichani et al., 2024a; Nguyen & Nguyen-Tang, 2025) that analyze the dynamics of attention logits in “discrete” settings where the attention logits diverge to infinity, the Coconut training method in continuous setting usually result in bounded attention logits. The bounded attention logits lead to a more smooth probability distribution over next tokens, which is beneficial especially under uncertainty: when the model is uncertain about the next step, a more smooth probability distribution under continuous CoT mechanism results in a superposition of different plausible next steps, which implements an effective exploration; on the contrary, an unbounded logit will result in a one-hot-like distribution and thus the model will over-confidently commit to a plausible branch and is likely to discard the ground-truth branch even when the evidence is weak.
@@ B.1 Analysis of Coconut-BFS
First, by Lemma 4, all $\dot{\mu}_{v}$ have the same value if all $\mu_{v}$ have the same value. Given that $\mu_{v}(0)=0$ for all $v\in\mathcal{V}$, we can obtain that for any time $t$, $\mu_{v}(t)$ have the same value for all $v\in\mathcal{V}$ using similar argument as in Lemma 15 of Huang et al. (2025).

### v3

@@ Abstract
Previous work shows that the chain of continuous thought (continuous CoT) improves the reasoning capability of large language models (LLMs) by enabling implicit parallel thinking, and a subsequent work provided theoretical insight by showing that a two-layer transformer equipped with continuous CoT can efficiently solve directed graph reachability by maintaining a superposition of multiple reasoning traces in the continuous thought. However, it remains unclear how the superposition mechanism is naturally learned from gradient-based training methods. To fill this gap, we theoretically analyze the training dynamics of a simplified two-layer transformer on the directed graph reachability problem to unveil how the superposition mechanism emerges during training in two training stages – (i) a thought-generation stage that autoregressively expands the continuous thought, and (ii) a prediction stage that converts the thought into the final answer. Our analysis reveals that during training using continuous thought, the index-matching logit, an important quantity which reflects the strength of the model’s local search ability, will first increase and then remain bounded under mild assumptions. The bounded index-matching logit effectively balances exploration and exploitation during the reasoning process: the model will exploit local problem structures to identify plausible search traces, and assign comparable weights to multiple such traces to explore when it is uncertain about which solution is correct, which results in superposition. Our experimental results tracking the growth of logits further validate our theory.
@@ 1 Introductions
Importantly, our analysis of the thought generation training stage reveals why the superposition can emerge even if the training data only presents one demonstration for each training sample. Our theoretical analysis as well as experimental results show that when training with continuous thought (i.e., the Coconut training method in Section 3 and Section 5), the index-matching logit, an important quantity that measures the strength of model’s local search capability, will remain bounded under mild conditions, which is in contrast to many previous analysis on transformer logit without continuous thought (e.g., Tian et al. (2023a); Nichani et al. (2024a); Nguyen and Nguyen-Tang (2025)) where the logit will grow logarithmically and thus unbounded. A bounded index-matching logit can balance exploration and exploitation: if the logit is too small, the model cannot even perform local search, resulting in a nearly random guess in the next step; if the logit is too large, the model might over-confidently commit to one of the plausible search traces merely depending on local features (e.g., the indegree of a node in our case) even if it is uncertain about the solution, and thus early discard the correct path. A bounded index-matching logit encourages the model to exploit the local structure while explore multiple plausible solutions by assigning comparable weights to them, which naturally results in superposition. This answers the question of Zhu et al. (2025) why superposition can emerge during training.
@@ 1.1 Related works
Chain-of-thought (CoT) (Wei et al., 2022) is a simple yet effective test time scaling method to enhance LLM’s reasoning capability. It can either be prompt-based only (Khot et al., 2022; Zhou et al., 2022) or be included in the training sample to create high-quality training data (Yue et al., 2023; Yu et al., 2023; Wang et al., 2023a; Shao et al., 2024). Beyond empirical study, many theoretical works also explore the advantages of the CoT method. For example, Liu et al. (2022); Feng et al. (2023); Merrill and Sabharwal (2023); Li et al. (2024b) shows that CoT can improve the expressivity of transformers. Zhu et al. (2024) studies the importance of CoT for two-hop reasoning via training dynamics. Wen et al. (2024); Kim and Suzuki (2024) studies how CoT in the training data can improve the sample efficiency of transformers. Instead of the text-based CoT, this paper studies continuous CoT where the “thinking tokens” lie in a latent continuous space and do not need to be converted to discrete tokens.
@@
A recent line of work studies latent space reasoning, a novel paradigm beyond text-based CoT (Goyal et al., 2023; Wang et al., 2023b; Pfau et al., 2024; Su et al., 2025; Hao et al., 2024). For example, Goyal et al. (2023) proposed to use pause tokens, which are learnable tokens that are inserted into the original text to increase the computation space. Later, London and Kanade (2025) theoretically shows that the pause token can strictly increase the expressivity of the transformer. Similarly, Pfau et al. (2024) studies filler tokens, which also increase the computation space of LLMs. Wang et al. (2023b) proposed to use planning tokens at the beginning of the response generation to improve the reasoning capability. Su et al. (2025) proposed to use abstract tokens in a latent space to enhance the reasoning performance while reducing the inference cost. The most related work is Hao et al. (2024), which proposes to use continuous CoT for reasoning. A follow-up work Zhu et al. (2025) theoretically shows the advantage of continuous CoT via expressivity. Our work takes a further step by analyzing the training dynamics of continuous CoT.
@@
There is a rich line of literature studying the optimization of transformer-based models (Jelassi et al., 2022; Bietti et al., 2023; Mahankali et al., 2023; Fu et al., 2023; Tian et al., 2023a; b; Zhang et al., 2024; Li et al., 2024a; Huang et al., 2024; Guo et al., 2024). More recent works also focus on understanding reasoning abilities or patterns of transformers via training dynamics, including induction heads (Nichani et al., 2024a), the reversal curse (Zhu et al., 2024; Ma et al., 2026), CoT (Wen et al., 2024; Kim and Suzuki, 2024; Huang et al., 2025b), factual recall (Nichani et al., 2024b), two hop reasoning (Guo et al., 2025; Lin et al., 2025), out of context reasoning (Huang et al., 2025a), etc. Along the line, our paper aims to understand the internal mechanism of continuous CoT and why superposition emerges via the analysis of training dynamics.
@@ 2 Problem Formulation
2.1 Index-matching Logits and Local Search Capability
Before we delve into the technical details in the following sections, we first provide an intuitive explanation of the dynamics of the main mechanism.
(a) Left: The continuous thought at step 1 [t1] encodes embeddings of nodes that are reachable from the root node r within one step. Middle: One-step expansion via local search where the strength is quantified by index-matching logit $\mu$. Right: After one-step expansion, the continuous thought at step 2 [t2] encodes nodes reachable within two steps.
(b) Illustration of how one-step expansion is implemented (adapted from Figure 3 of Zhu et al. (2025)). In the first layer of the transformer, each special edge token <e> copies its corresponding source and target nodes to its buffer spaces. In the second layer, as illustrated in the figure, the current thought [tc] pays large attention to an edge if its source node has been explored, and adds its target node to the superposition, where the strength of the added node is controlled by the index-matching logit $\mu$. The two edges $v_{5}\to v_{4}$ and $v_{1}\to v_{3}$ corresponds to edges in Figure 2(a).
Figure 2: Pictorial illustration of the superposition mechanism and the index-matching logit $\mu$.
Global planning vs. local search.
In the context of graph reachability, the global planning refers to a model’s capability to analyze the structure of the whole graph and then determine a path from the root node to the destination node. In contrast, local search focuses only on which nodes are reachable in one step from the current node, which is much easier to learn than global planning. When using discrete CoT, the model can choose only one path at a time. Therefore, the model needs global planning to select the correct path or to backtrack from the wrong one. When using continuous CoT, the model can keep multiple plausible paths simultaneously. Therefore, the model can rely solely on local search to perform parallel BFS, solving the task with only simple skills.
Index-matching logits.
[…]
@@ 3 Analysis of the Thought Generation Stage
Zhu et al. (2025) constructs a solution for a two-layer transformer, where the first layer mainly performs copy (e.g., the $i$-th special edge token <e> will copy the information of its corresponding source node $\texttt{s}_{i}$ and target node $\texttt{t}_{i}$). Since the copy mechanism has been widely studied (Nguyen and Nguyen-Tang, 2025), as well as its formation via training dynamics (Nichani et al., 2024a), we mainly focus on the dynamics after the copy mechanism has been established. Thus, we analyze the dynamics of the second layer of the transformer.
@@
where $\mathsf{E}_{\texttt{s}}(v)\in\mathbb{R}^{d}$ and $\mathsf{E}_{\texttt{t}}(v)\in\mathbb{R}^{d}$ map token $v\in\mathsf{Voc}$ to different subspaces of $\mathbb{R}^{d}$. For example, as in the construction of Zhu et al. (2025), we can set $d=3M$, and $\mathsf{E}_{\texttt{s}}(\cdot)$, $\mathsf{E}_{\texttt{t}}(\cdot)$ and $\mathsf{E}(\cdot)$ each corresponds to $M$ different non-zero entries. This is also similar to previous work Chen et al. (2025); Nguyen and Nguyen-Tang (2025) where $\mathsf{E}_{\texttt{s}}(\cdot)$ and $\mathsf{E}_{\texttt{t}}(\cdot)$ can be viewed as previous token heads. We make the following assumptions on the embedding $\mathsf{E}_{\texttt{s}}(\cdot)$, $\mathsf{E}_{\texttt{t}}(\cdot)$ and $\mathsf{E}(\cdot)$:
@@
where $\mathbf{V},\mathbf{W}\in\mathbb{R}^{d\times d}$ are attention parameters and $\sigma(\cdot):\mathbb{R}\to\mathbb{R}$ is an activation function that determines the range of attention scores, and $\bm{\xi}=(\xi_{v})_{v\in\mathsf{Voc}}\in\mathbb{R}^{M}$ is the output logit vector for each token in the vocabulary. Similar to the analysis in Nguyen and Nguyen-Tang (2025), we adopt the linear attention $\sigma({\bm{h}}^{\top}\mathbf{W}{\bm{h}}_{i})={\bm{h}}^{\top}\mathbf{W}{\bm{h}}_{i}$, fix $\mathbf{V}={\bm{I}}$ and use the index-matching reparameterization
@@
Note that, intuitively, the permutation-averaged loss will lead to similar behavior across different parameters. The first loss $\mathcal{L}^{\mathrm{BFS}}$ explicitly encourages the model to predict any nodes in $\mathcal{N}_{C+1}$. However, in practice, it is costly and even impossible to search over the entire solution space exhaustively; instead, we usually present only one demonstration for each task instance during training (in our setting, only one path $p$ per instance $(\mathcal{G},\texttt{r},\texttt{c}_{1},\texttt{c}_{2})$), which corresponds to the second loss $\mathcal{L}^{\mathrm{coco}}$ and aligns with the practical setting where chain of thought data can be used for supervision.
@@
By Theorem 1, as long as $F(0)<d_{p_{c+1}}<d_{\max}$, we have $\mu(t)\to\mu^{*}>0$. Compared to many previous work (Tian et al., 2023a; Nichani et al., 2024a; Nguyen and Nguyen-Tang, 2025) that analyze the dynamics of attention logits in “discrete” settings where the attention logits diverge to infinity, the Coconut training method in continuous setting usually result in bounded attention logits. The bounded attention logits lead to a more smooth probability distribution over next tokens, which is beneficial especially under uncertainty: when the model is uncertain about the next step, a more smooth probability distribution under continuous CoT mechanism results in a superposition of different plausible next steps, which implements an effective exploration; on the contrary, an unbounded logit will result in a one-hot-like distribution and thus the model will over-confidently commit to a plausible branch and is likely to discard the ground-truth branch even when the evidence is weak.
@@ B.1 Analysis of Coconut-BFS
First, by Lemma 4, all $\dot{\mu}_{v}$ have the same value if all $\mu_{v}$ have the same value. Given that $\mu_{v}(0)=0$ for all $v\in\mathcal{V}$, we can obtain that for any time $t$, $\mu_{v}(t)$ have the same value for all $v\in\mathcal{V}$ using similar argument as in Lemma 15 of Huang et al. (2025a).
@@ E.3 Alternative Attention Routes for Candidate Lift
Appendix F Additional Experimental Results
In this section, we provide additional experiments to complement our main analysis of training dynamics.
F.1 Ablation Study: Architectural and Optimization Sensitivity
We evaluate the sensitivity of Coconut training to model depth, number of attention heads, hidden width, and learning rate. The results are summarized in Table 4.
Number of layers
$L=2$ $98.8$
$L=4$ $97.3$
$L=8$ $96.5$
$L=12$ $67.4$
Number of heads
$H=4$ $98.0$
$H=8$ $98.8$
$H=12$ $98.8$
Width
$d_{\text{model}}=384$ $62.0$
$d_{\text{model}}=768$ $98.8$
$d_{\text{model}}=1536$ $97.7$
Learning rate
$\eta=2\times 10^{-4}$ $58.1$
$\eta=1\times 10^{-4}$ $98.8$
$\eta=5\times 10^{-5}$ $62.1$
Weight tying
Tied $98.8$
Untied $98.8$
Table 4: Ablation on depth, heads, width, learning rate, and weight tying. By default, other hyperparameters follow the main experiments.
We observe that models with $L=\{4,8\}$ layers maintain high accuracy, while $L=12$ is harder to optimize. The performance remains comparable when $d_{\text{model}}\in\{768,1536\}$, but degrades when the width is too small (e.g., $d_{\text{model}}=384$). Varying the number of heads does not have major effects on final accuracy, whereas too large or too small learning rates tend to degrade performance. Weight tying setting does not affect model performance.
We emphasize that each ablation in Table 4 varies only a single hyperparameter at a time, keeping all other settings identical to our main experiment. In practice, these hyperparameters interact in a coupled manner. For instance, with a smaller learning rate of $5\times 10^{-5}$, we can extend the first-stage training to 300 epochs and get 97.0% accuracy. For deeper models with $L=12$, prolonging first-stage training to 400 epochs and reducing the learning rate to $5\times 10^{-5}$ improves accuracy to 99.6%. A comprehensive hyperparameter interaction study is beyond the scope of this work and is left for future investigation.
F.2 Multi-Layer Transformers and Mechanistic Patterns
We use the Coconut model with $L=4$ to analyze the reasoning pattern beyond two-layer transformers. The results are shown in Figure 7 and Figure 8, and we summarize the reasoning patterns below.
  • First layer (induction head): The first layer performs token-level copying, propagating node information into edge tokens <e>, consistent with the copy mechanism derived in previous theoretical analysis (Zhu et al., 2025).
  • Second layer and beyond (superposition): From the second layer onward, the model aggregates over reachable nodes in a superpositional representation that enables parallel breadth-first exploration.
Figure 7: The first-layer attention patterns in 4-layer transformers. <e> tokens attend to the corresponding source and target nodes to aggregate the information. This is consistent with the analysis of the two-layer transformer in Zhu et al. (2025).
Figure 8: Inner product between layer-wise hidden states and different types of nodes in a 4-layer transformer. The experimental setting follows Zhu et al. (2025). From the second layer onward, hidden states exhibit larger inner products with reachable, frontier, and optimal nodes, indicating that superpositional representations emerge as early as layer 2 in the 4-layer transformer.
F.3 Accuracy Dynamics in the Answer-Prediction Stage
We track the test accuracy during the final answer-prediction stage following the setting in Figure 4. The result is shown in Figure 9, which shows a rapid transition from near-random guessing to stable high accuracy once the model integrates residual carryover and candidate lift signals.
Figure 9: Accuracy curve during the answer-prediction stage. The accuracy shows a rapid improvement corresponding to the learning of residual carryover and candidate lift signals.
Appendix G The Use of Large Language Models (LLMs)
We used LLMs mainly for grammar checking and polishing in paper writing.